OpenStudy (anonymous):

use abslon-delta to show that f(x)=x^2 AT x dote =2

OpenStudy (anonymous):

Abslon!

OpenStudy (anonymous):

what is \(x\) approaching?

OpenStudy (anonymous):

yes abslon

OpenStudy (zzr0ck3r):

epsilon :)

OpenStudy (zzr0ck3r):

what is dote?

OpenStudy (anonymous):

xo

OpenStudy (zzr0ck3r):

what is xo ?

OpenStudy (zzr0ck3r):

\[\lim_{x\rightarrow?}x^2\]

OpenStudy (anonymous):

its continuous property

OpenStudy (anonymous):

when we use epsilon - delta and say x^2-2 <EPSLION

OpenStudy (anonymous):

Does x approach \(2\)? Does x approach \(\sqrt 2\)?

OpenStudy (anonymous):

the first on x approach 2

OpenStudy (anonymous):

f(x)=x^2 at x0 =2

OpenStudy (anonymous):

Well as \(x\to 2\) then \(x^2\to 4\). Right?

OpenStudy (anonymous):

\[ |x-2|<\delta \implies |x^2-4|<\epsilon \]

OpenStudy (anonymous):

\[ |x-2|<\delta \implies |x-2||x+2|<\epsilon \]

OpenStudy (anonymous):

I have to say first and prove /x^2-2/epsilon

OpenStudy (anonymous):

< epsilon

OpenStudy (anonymous):

Okay sorry, but you obviously are confused..

OpenStudy (anonymous):

\[ \lim_{x\to a}f(x)=L \]You need to decide what \(a\) and \(L\) are. Let me assure you they both can't be \(2\) because \(2^2\neq 2\)

OpenStudy (anonymous):

yes sorry I know because its hard to write the symbol

OpenStudy (anonymous):

yes I think I have to prove that f(x) is continues if f(x)=x^2 at x0= 2

OpenStudy (zzr0ck3r):

to denote \(x_0\) write x_0

OpenStudy (zzr0ck3r):

ok do you need to prove something is continuous or do you need to prove the limit?

OpenStudy (anonymous):

okay sorry about yhat

OpenStudy (zzr0ck3r):

no need to be sorry:)

OpenStudy (anonymous):

I need to prove it continues

OpenStudy (zzr0ck3r):

so you need to prove x^2 is continuous at x = 2?

OpenStudy (anonymous):

RIGHT

OpenStudy (anonymous):

So show \(f(2) = L\) and then show \(\lim_{x\to 2}f(x) = L\)

OpenStudy (anonymous):

No abslons needed.

OpenStudy (anonymous):

no in the question say use epsilon - delta

OpenStudy (anonymous):

Okay so first let's find out \(L\). Can you do that?

OpenStudy (perl):